OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x^2 + 3x + 2 y = x^2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x^2 + 3x + 2 different from y = x^2 + 2x + 1?

OpenStudy (earthcitizen):

All that can be said to this q is Wolfram is your friend, First, y =(x+1) (x+2), x=-1 , -2 Second, y = (x+1)^2 = 0, x = -1

OpenStudy (campbell_st):

y = x^2 + 3x + 2 factorised to y =(x+2)(x +1) the parabola is concave up, intercepts the x axis at -2, -1 intercepts the y axis at + 2 has a line of symmetry of x = - 1.5 has a minimum value of y = -1/4 the 2nd curve is y = x^2 + 2x + 1 factorised is y = (x +1)^2 the parabola is concave up has 1 point of contact with the x axis at x = -1 has y intercept at y = 1 the has a lie of symmetry at x = -1 has a minimum value of y = 0 the commonality in the graphs is that they are both concave up have a common x intercept of x = -1